3/4a + 1=2/3a - 1/4a
Answer by checkley79(3341)

3/4a + 1=2/3a - 1/4a
3A/4+1=2A/3-A/4
3A/4-2A/3+A/4=-1
(3A*3-2A*4+A*3)/(3*4)=-1
(9A-8A+3A)/12=-1
4A/12=-1
4A=-12
A=-12/4
A=-3 ANS.
PROOF:
3*-3/4+1=2*-3/3-(-3)/4
-9/4+1=-6/3+3/4
-2.25+1=-2+3/4
-1.25=-2+.75
-1.25=-1.25